### v2

# Template-Based Visual Program Distillation

Abstract
For users with limited computational resources, visual programming or prompting large language models (LLMs) to generate executable code for visual tasks, like visual question answering (VQA), remains largely inaccessible. Even with techniques such as distillation, adapting visual programming to smaller models or specific datasets is still quite challenging due to high annotation costs. We propose a low-cost visual program distillation method that can be used for models with fewer than 1 billion parameters and requires no human-generated program annotations. We achieve this through synthetic data augmentation based on decoupling programs into higher-level skills, called templates, and their corresponding arguments. Experimental results show that, with a relatively small amount of question/answer data, small language models can generate high-quality visual programs with the added benefit of much faster inference.
Template-Based Visual Program Distillation
Michal Shlapentokh-Rothman  Yu-Xiong Wang  Derek Hoiem University of Illinois at Urbana-Champaign {michal5,yxw,dhoiem}@illinois.edu
1 Introduction
Visual programming Gupta and Kembhavi (2023); Subramanian et al. (2023); Surís et al. (2023) refers to generating programs that invoke visual models to solve tasks such as answering questions about images, typically by prompting a very large language model (LLM) like GPT Achiam et al. (2023) or Llama Touvron et al. (2023). Despite success in zero-shot settings, visual programming remains largely inaccessible due to high program generation costs, which come from the need for expensive hardware to run LLMs or using expensive proprietary APIs, as well as long inference time and high data costs for both human-generated program annotations and question/answer annotations. Previous efforts Khan et al. (2024) around generating visual programs on smaller LLMs have made some progress in addressing these concerns, but training and inference costs remain substantial.
Our goal is to create a visual programming system with two key characteristics:
  1. 1.
    Small Program Generator Programs should be generated by models with $<1$ billion parameters, enabling fast inference and use on a wide variety of hardware.
  2. 2.
    Limited Human Annotations Minimize the number of required annotations, especially human-generated programs.
To achieve this, our approach builds on the key observation that many visual programs share a common high-level structure – a unique aspect of visual programming. For example, the programs ‘Count the red chairs’ and ‘Count the green apples’ follow the same pattern but differ in their inputs. Inspired by this insight and recent advances in LLM distillation Hsieh et al. (2023), we propose a low-cost visual program distillation method. Our approach requires no human-generated example programs and uses only a small fraction of question/answer pairs (at most 1000). A teacher model leverages auto-context generation, where each generated program that produces the correct answer is added to the set of in-context examples. The resulting annotated programs form a dataset used to train a small language model (SLM) to generate visual programs efficiently.
Method Task Approach Model Size Train Size Domain
VisRep Khan et al. (2024) Program Gen. Self-Training 7B 10% ($\approx$10K) Visual Prog.
Distill Step-by-Step Hsieh et al. (2023) Text Gen. LLM Distillation 220M-770M 12.5% Q/A, Reasoning
CodePlan Sun et al. (2024) Program Gen. LLM Distillation 770M 100% General Coding
Template-Based Aug. (Ours) Program Gen. LLM Distillation 770M 0.1% ($\approx$1K) Visual Prog.
Table 1: Comparison of approaches for program generation and LLM distillation. Our method enables LLM distillation for visual programming using only a small fraction of the training dataset and minimizing annotation costs.
During training, program annotations are synthetically generated through template-based augmentation. This strategy decouples high-level skills or plans, called templates, from question-specific concepts or arguments. For example, the template for ‘Count the red chairs’ is ‘find(arg1), verify property(arg2), count’ and the corresponding arguments are ‘red’ (arg1) and ‘chair’ (arg2). New question/program pairs are created by replacing these arguments in the question and program. By replacing ‘red’ with ‘green’ and ‘chairs’ with ‘apples,’ we generate both a new question and a corresponding program. Compared to directly training on the program annotations, training with template-based augmentation offers three key benefits: 1) better performance when less data is available, 2) increase in student/teacher prediction agreement, and 3) higher-quality generated programs. Our method uses a low-cost teacher model (e.g. GPT-Mini-4o) and costs only around \$1 per dataset. In addition, the distilled models achieve much faster inference speed than existing LLMs.
To validate our approach, we evaluate on frequently used visual question answering (VQA) datasets and compare with using human-generated annotations as well as non-augmented distillation. Moreover, we address an overlooked evaluation aspect by performing several types of evaluation, including human program verification, due to the noisiness of commonly used metrics.
All together, our results show that it is possible with few question/answer pairs to distill the visual program ability of LLMs into a much smaller model that does not require in-context examples.
2 Related Works
We outline the most relevant related work below. A high-level comparison between our work and others can be found in Table 1.
Visual Programming
A long line of work investigates generating and executing programs to perform visual tasks. Early approaches (Andreas et al., 2015, 2016; Hu et al., 2017; Johnson et al., 2017) generate programs and execute the programs with learned end-to-end neural modules. Based on the impressive code generation capabilities of LLMs (Bareiss et al., 2022), visual programming frameworks (Surís et al., 2023; Gupta and Kembhavi, 2023; Subramanian et al., 2023) generate programs given an API and in-context examples and execute the programs with large pre-trained vision models. Visual programming has been applied to many different domains and applications including visual question answering, video question answering, text-to-image generation, and robotics (Cho et al., 2024; Lu et al., 2024; Min et al., 2024; Liang et al., 2023). Unlike previous visual programming works, we only use small models ($<1$ billion parameters) to generate programs.
Visual programming is different from both LLM tool use and LLM code generation since it requires both basic reasoning and knowledge skills as well as basic code generation. One of the challenging parts of visual programming comes from the execution and evaluation due to underlying models. As noted in other works Khan et al. (2024), incorrect programs can return (after being executed) the correct answer even if the program is incorrect.
Tool-Based Finetuning
Compared to prompt-based methods, there are relatively few methods focused on improving program generation in visual programming or in the general field of tool-based LLMs through finetuning. One of the main challenges is the lack of program annotations for input/output pairs. Language modeling has a long history designing self-supervised tasks, especially in pre-training (Devlin et al., 2019; Raffel et al., 2020). Toolformer (Schick et al., 2024), uses a form of self-supervision, to create a training dataset to finetune an LLM on tool-use programs. For each question/answer pair in a pre-existing training dataset, an LLM is prompted to generate a corresponding program. If the program decreases the training loss (of the same LLM) then the program annotation is added to a new dataset. Then the same LLM is finetuned on the generated dataset. A similar approach is used in Chain-of-Thought (CoT) finetuning (Zhu et al., 2023), program correction methods such as VDebugger (Wu et al., 2024) and visual programming based finetuning methods (Khan et al., 2024) such as VisRep. LoRA (Hu et al., 2021) is frequently used when finetuning LLMs on the generated data.
In VisRep, Khan et al. (2024) use a self-training approach similar to toolformer for visual programming. Self-annotated programs are kept if the executed program returns the correct answer. While there are some similarities between VisRep and our work, the underlying setting is different. VisRep focuses on improving program generation on an existing LLM (7B) through self-training, while the goal of our work is to distill visual program generation from an LLM to a small model (<1B). Our work also does not require sampling across specific question types or correcting by hand different program annotations during training. We randomly sample the training set and do not write any program annotations or corrections.
Distillation
Knowledge distillation (Hinton et al., 2015; Buciluǎ et al., 2006), where a large teacher model annotates unlabeled data that is then used to finetune a smaller, weaker student model, is frequently used across many different applications including image recognition (Beyer et al., 2021), masked language modeling (Sanh et al., 2019) and commonsense knowledge (West et al., 2021). One of the difficulties of distillation is the need for a large number of unlabeled training examples which can be expensive to obtain. One way to compensate for this is to train the student model with a multi-task objective. The objective consists of a weighted sum of the cross-entropy on the original task and cross-entropy on a closely related part of the task such as the rationale in CoT (Hsieh et al., 2023). CodePlan (Sun et al., 2024) applies such an idea to code generation where the secondary objective is a natural language version or plan of the code, which has been shown to be effective in prompting-based work (Jiang et al., 2024).
One downside to multi-objective prompting is that both the teacher and the student models have to generate additional output. In CodePlan, during inference, the model first generates the plan and then generates the desired code creating long inference time. Such steps might be necessary for complex code generation but are unnecessary for visual programming. Instead we use a relatively simple data augmentation method based on abstractions (Yuan et al., 2024) or higher level plans of existing programs which we refer to as templates. Templates do not require any additional forward passes to create and can be augmented using simple techniques such as word replacement Tang et al. (2019).
Automatic In-Context Example Generation
Prompt-based methods are more effective when in-context examples or demonstrations (Wei et al., 2022) are included. In visual programming (Surís et al., 2023; Gupta and Kembhavi, 2023), manually written in-context examples are used to adapt an API to a particular dataset. However, such an approach can be time consuming and not feasible for a large training set of input/output pairs. A common practice is to have an LLM self-annotate examples and keep the examples that produce the correct answer Zhang et al. (2023); Stanić et al. (2024); Tao et al. (2024); SU et al. (2023). Then during inference, augmented generation (RAG) Lewis et al. (2020)can be performed to select the most relevant ones. We follow such an approach, referred to as auto-context generation by  Tao et al. (2024) to generate in-context examples for the teacher model.
Template image_patch = ImagePatch(image)
var1 = image_patch.find(‘arg_0’)
var2 = var2.classify(‘arg_1’)
var3 = image_patch.find(‘arg_2’)
var4 = var3.classify(‘arg_3’)
answer = bool_to_yesno(var2 == var4)
Are the cat and the tshirt the same color? arg_0 = cat arg_1 = color arg_2 = tshirt arg_3 = color
Is the sofa made from the same material as the chair? arg_0 = sofa arg_1 = material arg_2 = chair arg_3 = material
Is the vase the same shape as the table? arg_0 = vase arg_1 = shape arg_2 = table arg_3 = shape
Table 2: A template is a particular ordering of operations. The questions above all share the same template since they only differ in the arguments. We want to answer similar questions the same way and easily generate synthetic data.
3 Method
Preliminaries
Following the notation in ViperGPT Surís et al. (2023), the visual programming objective is to generate a program $z=\pi(q,p)$ with a program generator, $\pi$, input query $q$ and prompt $p$ such that when executed with execution engine $\phi$ and corresponding visual input $x$, $\phi(x,z)$ returns the correct answer. Prompt $p$ contains an API and dataset-specific in-context examples.
Figure 1: An overview of our augmentation method. Programs are first separated into templates and argument, new arguments are selected and plugged back into the question/program pair. Templates are created by renaming variables and removing question specific concepts. One single teacher generated question/program pair can turn into hundreds of new question/program pairs.
Given a teacher model, $\pi_{t}$, and a smaller student model ($<1$ billion parameters), $\pi_{s}$, our goal is to distill visual program generation from the teacher to the student for a specific dataset using a minimal number of answer annotations (question/answer pairs). We do not have any human provided annotations (question/program/answer triplets) but we do know the API used for the teacher model.
There are three main steps to our approach: teacher annotation, data augmentation and student training.
3.1 Teacher Generated Program Annotation
We use GPT-4o-Mini AI (2024) as our teacher model. Unlike traditional knowledge distillation, LLM based distillation requires the teacher (an LLM) to be given the appropriate context through in-context examples before starting annotation. Common practice is to have a human generate the examples. We follow a process known as Auto-Context Generation to automatically generate such examples given answer annotations only.
Figure 2: An example of our data augmentation approach. Both the new and old question have the same template so the template matcher output should predict the same template for both. The arguments for the new and old programs are different. But, in the arguments, (dog, sofa, brown) should be replaced with (bear, desk, green).
Auto-context generation follows a simple process similar to VisRep Khan et al. (2024):
  1. 1.
    Teacher model predicts a program using API and in-context examples (if any).
  2. 2.
    Generated program is evaluated.
  3. 3.
    If the answer returned by a generated program matches the ground truth answer, then the program is immediately added to the set of in-context examples. Otherwise the program is discarded.
For efficiency, the in-context examples are sampled based on similarity to the question (i.e. RAG) Lewis et al. (2020) once there are more than 50. We compute the cosine similarity with finetuned MP-Net-v2 Reimers and Gurevych (2019) between an input and all in-context examples and select the 50 highest scoring examples.
3.2 Data Augmentation
After the teacher generates a dataset of question/program pairs, our goal is to train a student model on these pairs only. At this point, corresponding answers and images of questions are not used. Since the dataset is small, we use data augmentation to create a greater variety of question/program pairs.
To understand the intuition behind our data augmentation method, consider the set of questions in  Table 2. All of these questions compare properties of two objects. The general structure of each program is the same with the only difference coming from the inputs to the functions. If we answer one of these questions correctly and know that the remaining questions have the same structure, then all of the remaining questions should have that same structure or should be consistent. An overview of our data augmentation method can be seen in Figure 1. We refer to this as ‘template-based augmentation.’
Templates
We define a template as a specific ordering of functions, where a function is an API call to a visual model or python operation. Templates are argument independent. For example, if the program is
image_patch = ImagePatch(image)
dog = image_patch.find(‘dog’)
answer = dog.classify(‘color’)
then the template would be
image_patch = ImagePatch(image)
var1 = image_patch.find(<arg>)
answer = var1.classify(<arg>)
Please see Table 3 for examples of questions and corresponding templates.
Templates can be considered high-level plans used for plan-based distillation Sun et al. (2024). The main advantage to templates is that they can be extracted directly from the program. There is no need to generate extra output or perform multiple forward passes for a single question.
Template Extraction
Given the program annotations from the teacher model, we extract templates and corresponding arguments from each program similar to the abstraction method of CRAFT Yuan et al. (2024). Extracting is quite simple: replace specific variable names with generic ones and put in placeholders for each argument. The variable renaming and extraction can be done in seconds with abstract syntax trees and regular expression matching. The code can be found in Appendix E.
Augmentation
We can use the decomposition of programs into templates and arguments to generate synthetic data similar to how masked language modeling is used in BERT Devlin et al. (2019). As can be seen in Table 2, for many questions and programs, the arguments appear directly in the question. Consider the example in Figure 2: “Are both the dog and the sofa brown?" The arguments are dog, sofa, and brown. Once we find similar words for each, we can simply replace them in the sentence. Since we already know the template and the arguments, we also have the program. For GQA (Hudson and Manning, 2019), the possible word replacements are from ViperGPT Surís et al. (2023) and for arguments that are either not in the question or are equal to the entire question, we use GPT to generate similar questions. For VQAv2 Goyal et al. (2017), we use BERT Devlin et al. (2019) when an argument is a single word and BART Lewis (2019) to replace phrases. If possible, we replace arguments with arguments of the same type e.g. an attribute is replaced by an attribute.
Example Template
Is the blue car the same shape as the chair? Is leather jacket made of the same material as the shirt? image_patch = ImagePatch(image)
var1 = image_patch.find(<arg_0>)
var2 = filter_img(<arg_1>)
var3 = var2.classify(<arg_2>)
var4 = image_patch.find(<arg_3>)
var5 = var4.classify(<arg_4>)
answer =bool_to_yesno(var3 ==absent$=$ = var5)
What type of food is near the person? What is the vehicle next to the animal? image_patch = ImagePatch(image)
var1 = image_patch.find(<arg_0>)
var2 = var1.crop_position(<arg_1>)
var3 = var2.find(<arg_2>)
answer $=$ var3.classify(<arg_3>)
Is the car to the left or right of the tree? Is pot above or below the pan? image_patch = ImagePatch(image)
var1 = image_patch.find(<arg_0>)
var2 = image_patch.find(<arg_1>)
answer $=$ choose_relationship(var1,var2,<arg_2>)
Table 3: Some examples of questions and corresponding templates. Multi-colored words correspond to multiple arguments.
During training, each argument in each question has a 50% chance of being replaced. If an argument is to be replaced, we then uniformly sample among the possible replacements. In the GQA dataset, for some arguments, like objects, the number of possible replacements is quite large (e.g. greater than 1500), while for arguments like directions such as left, right, etc. the number of replacements is small (e.g. fewer than 10). Some examples of categories and more details on word replacement can be found in Section F.
Student Training
Given the augmented dataset, we perform LoRA-based finetuning on the student model. Since the augmentation method produces both questions and corresponding programs, we use next-token prediction loss for training.
4 Experiments
Experimental Setup
For all experiments, the teacher model is GPT-4o-mini, the student model is CodeT5 Wang et al. (2023) and is trained with LoRA Hu et al. (2021). We use the GQA dataset Hudson and Manning (2019) and VQAv2 Goyal et al. (2017) datsets for our experiments. We evaluate on the full GQA test-dev split and randomly sample 10000 questions from the VQAv2 validation split. GQA is evaluated using exact match and VQAv2 is evaluated based on annotator/answer agreement as in the original benchmark.
We use a slightly modified API from the original ViperGPT paper  Surís et al. (2023). The main differences are some additional functions (to reduce program length) and removal of the use of a VLM if earlier parts of a program fail. Please see Appendix Section C for more details. For visual models, we use InstructBLIP (Flan-T5 XL) Dai et al. (2023) for general visual queries, Owl-ViT2 Minderer et al. (2024) for detection and CLIP Radford et al. (2021) for classification.
Auto-Context Generation
Before training the student model, we adapt the teacher to the VQA domain. In Table 4, we investigate the effectiveness of using human-generated programs and programs generated by the teacher model validated by answer correctness. As our aim is to enable visual programming with low cost and effort, we compare using a small number (25) of human-generated programs to generating 1000 programs (yielding hundreds of validated programs).
Dataset Human Generated Auto-Context Generated Performance
GQA 0 0 35.1
25 0 38.5
0 474 (out of 1000) 43.1
25 510 (out of 1000) 43.1
VQAv2 0 0 47.4
25 0 54.2
25 283 (out of 500) 57.8
0 286 (out of 500) 60.3
Table 4: Teacher performance compared across varying numbers of human-generated and auto-context-generated in-context examples.
As shown in Table 4, human-generated programs provide significant improvement vs. no examples, but provide no further benefit given self-generated examples. This finding supports using self-generation to improve the teacher model, as question-answer pairs require much less expertise and time to provide.
Data Usage GQA VQAv2
Aug. No Aug. Aug. No Aug
0.1% 41.9 (+0.8) 41.1 61.1 (+0.3) 60.8
0.05% 39.7 (+4.1) 35.6 59.9 (+9.1) 50.8
0.02% 34.2 (+4.2) 30.0 49.0 (+7.8) 41.2
Table 5: Answer accuracy with and without augmentation across different dataset sizes: 1000 (0.1%), 500 (0.05%), and 250 (0.02%) for GQA and 500 (0.1%), 250 (0.05%) and 125 (0.02%) for VQAv2. Performance gain from augmentation is higher when using 0.05% and 0.02% of the dataset compared to 0.1%.
Augmentation vs Non-Augmentation
When training the student model, we use templates and variable substitution to create a more diverse set of programs for training. We evaluate the effect of template-based data augmentation in two ways. In Table 5, we analyze answer accuracy as a function of varying dataset sizes. The results indicate that data-augmentation increases the answer accuracy of the student model, with the largest benefits when few generated programs are available for training.
To better evaluate the effectiveness of distillation, we also evaluate whether the teacher and student provide the same answer. In Table 6, we see that augmentation leads to much higher rates in student-teacher agreement, even when the overall improvement in accuracy is small (see Table 5).
GQA VQA
With Augmentation 64.5 76.1
No Augmentation 59.8 72.0
Table 6: Student/Teacher Prediction Agreement. Data augmentation has a much larger effect on student/teacher prediction compared to answer accuracy.
Cost and Efficiency
One of our objectives is keeping annotation costs low. As shown in Table 7, the total annotation cost for program auto-generation on both datasets is less than a dollar. GQA costs a bit more because of the larger size and during augmentation, GPT-4o-Mini is called when arguments do not appear in the question. By using a small model, inference speed also greatly increases.
GQA VQAv2
Annotation Cost \$0.69 \$0.26
CodeT5 Inference Time (q/s) 35 32.7
Llama-3.2-1B Inference Time (q/s) 0.32 0.33
GPT-4o Mini Inference Time (q/s) 1.3 1.3
Table 7: Annotation cost and inference time for several different models. Our method has low cost and high inference speed with CodeT5.
Comparison with Related Work and Program Evaluation
We compare our approach with the most closely related setting: prompt-based visual program generation. Other works, such as VisRep Khan et al. (2024), have different objectives and use carefully curated and larger datasets, and thus are not included in our comparisons. Furthermore, VisRep evaluates on selected subsets of various datasets, and the specific subsets, models, and code have not been made publicly available for comparison.
Method GQA VQA
Answer Acc Program Acc Answer Acc Program Acc
GPT-4o-Mini (w/Human Examples) 38.5 78 54.2 92
Llama 3.2 1B (w/Human Examples) 19.0 22 23.3 40
Code-T5 Distilled w/o Augmentation (Ours) 41.1 48 60.8 79
Code-T5 Distilled w/Augmentation (Ours) 41.9 68 61.1 80
GPT-4o-Mini (Auto-context) (Ours) 43.1 85 60.3 93
Table 8: Answer and program accuracy of different methods on GQA and VQAv2. With augmentation, distilled programs can achieve similar or better answer accuracy as prompting methods and improve program accuracy compared to not using augmentation.
The widely followed ViperGPT Surís et al. (2023) approach is to use a prompt containing an API and a set of human generated in-context examples. We use two models to evaluate prompting GPT-4o-Mini and Llama-3.2-1B Dubey et al. (2024), a distilled version of Llama 3. 25 human generated in-context examples are used per dataset for each model. In Table 8, we refer to these baselines as GPT-4o-Mini (w/human examples) and Llama-3.2-1B (w/human examples). From our method, we evaluate on two distilled CodeT5 models with and without augmentation. The distilled models are trained with 0.1% of the data (474 question/answer pairs for GQA and 286 question/answer pairs for VQAv2). We also include the auto-context teacher model, notated as GPT-4o-Mini (Auto-context).
Table 9: Common errors and examples used during qualitative evaluation.
Error Question Program Explanation
Not executable Is the chair on the left or right side? chair = image_patch.find(‘chair’) side = choose_relationship(chair, image_patch, ‘left or right’) choose_relationship requires a list as input
API Violation What color is the running dog on the left? left = crop_position(‘running left’, image_patch) dog = left.find(‘dog’) color = dog.classify(‘color’) ‘Running left’ is not a direction or preposition
Contradicts Question What color is the car above the road? road = image_patch.find(‘road’) below_road = crop_position(‘below’, road) car = below_road.find(‘car’) car_color = car.classify(‘color’) The question states that the car is above the road, not below.
Does Not Answer Question Are there two tables? tables = image_patch.find(‘table’) num_tables = count(tables) The question asks if there are two tables, not how many tables there are.
Does not include all question information Is the blue toy small? toy = image_patch.find(‘toy’) is_small = toy.verify_property(‘small’) answer = is_small The question specifically asks about the blue toy.
We evaluate using two metrics: answer accuracy and program accuracy. Program accuracy involves a human (the authors) manually evaluating each program (without execution) for correctness. We randomly sample 100 questions from each dataset for evaluation. The same 100 questions are used to evaluate each method.
There are multiple ways to determine if a program is correct. We generally assume a program is correct unless it violates one of these criteria:
  • Not Executable: The program must be executable and return the correct data type (a string for VQA datasets).
  • API Violation: Visual programming APIs are designed to follow basic visual knowledge and reasoning. For example, the ‘find’ function is used with nouns while ‘verify_property’ is generally used for attributes. Clear violations such as trying to find an attribute (e.g. ‘find(green)’) or cropping with a verb (e.g. ‘crop_position(running)’) are considered incorrect.
  • Contradicts Question: Programs that assume a statement that directly conflicts with a statement in the question. For example, assuming an object is on the left, when the question states it is on the right.
  • Does Not Answer Question Programs that do not answer the question, even if the program correctly follows the API, are incorrect. Common examples are returning yes/no instead of choosing between two options such as left or right.
  • Does Not Include Vital Information From the Question If the question includes details about an object, then those details must be in the program.
Some examples can be seen in Table 9.
We can draw several observations from the results in Table 8. From the answer accuracy results, we see that with a small amount of data and a strong enough teacher, student models can achieve similar performance to the auto-context trained teacher. Auto-context generation also improves the teacher performance. The performance from prompting Llama is particularly low, consistent with previous findings showing that it is difficult to transfer prompts across LLMs Sclar et al. (2024) of similar size, let alone much smaller ones.
The program accuracy results are bit more surprising. For all of the methods, program accuracy is higher than answer accuracy especially for GPT-4o-Mini (Auto-context) which has 85% vs 43% and 93% vs 60% accuracy for GQA and VQAv2 (respectively). Data augmentation has a larger effect on program accuracy than question accuracy but a substantial gap still remains between the student and teacher models, indicating more room for improvement. Across the two datasets, we see a much higher overall program accuracy for VQAv2 than for GQA. Both datasets have fairly easy questions but VQAv2 has many that are correctly answered by either calling a vision-language model (the simple_query) function or simple counting, which involves two functions (find and count). Taken together, our evaluation suggests that the quality of the visual programs is much better than what answer accuracy alone would imply. We include additional analysis of program errors in Appendix A.
5 Conclusion
Our experiments show with template-based visual program distillation, only a small number of answer annotations are needed to achieve similar performance to prompt-based methods. Auto-context generation removes the burden of human generated program annotations while still retaining the same performance. Template-based augmentation improves both student/teacher prediction agreement and program accuracy compared to non-augmented approaches. The distilled student models achieve similar answer accuracy as the teacher at a fraction of the time and cost.
In conclusion, we show that the visual program ability of LLMs can be distilled into much smaller models with only a few question/answer pairs and no human-generated programs. For less than \$6, (\$5 for student training on the cloud and \$1 for annotation), a 770-M coding model can become a visual program generator. We anticipate that use of template-based visual program distillation will enable users and researchers to iterate more quickly on various visual programming systems and broaden their applications.
6 Limitations
There are several limitations to our work:
Indirect Evaluation Answer accuracy is an indirect evaluation of program correctness. As indicated by our program accuracy results shown in Section 4, answer-based evaluation is only a loose indicator of visual program correctness. Promising areas for future work include automatic program accuracy evaluation and closing the gap between answer accuracy and program accuracy.
Reliance on teacher model and API Our method relies on the quality of the teacher model, which in turn is dependent on the quality of the API and general prompt. A future challenge is how to learn from a weaker teacher and/or unreliable API.
Teacher Data Efficiency Following prior works, if a program was incorrect, it was not used. However, LLMs have the ability to self-correct given the appropriate feedback. In settings where question/answer pairs are limited, such an approach could be more cost-effective than discarding the examples. Areas for future work include incorporating self-correction methods or using program correction models such as VDebugger Wu et al. (2024).
Program Execution Time Our method decreases program generation time, but not program execution time. With a single GPU, program execution time on average was 3.37 s and varies quite widely. The long execution time remains a bottleneck in different applications of visual programming.
Limited Program Complexity Existing VQA datasets are fairly simple and most work on visual programming is limited to tasks where the programs can be generated in a single step by an LLM. Most real world applications are multi-step and would require more complex reasoning and knowledge skills than in the evaluation datasets.
Acknowledgement
This research is supported in part by NSF Grant 2106825, NIFA Award 2020-67021-32799, ONR N00014-23-1-2383 and ONR N00014-21-1-2705. This work used the Delta system at the National Center for Supercomputing Applications through allocation CIS230398 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program Boerner et al. (2023), which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix A Program Analysis
Program Errors
In Figure 3, we show the frequency of different error types for each method on the GQA dataset. Note that incorrect programs can fall into multiple categories but in this classification, each incorrect program was counted only once. There are very few execution errors and most errors come from either contradicting the question or not answering it. Many of the Llama generated programs returned yes/no even if the question asked for a different string. Augmentation largely reduces such errors but also generates new errors.
Figure 3: The frequency of errors across the different categories for GQA program evaluation. Augmentation reduces the number of ‘Does Not Answer Question Mistakes.’
Qualitative Analysis
In Figure 4, we show 9 generated programs from 3 questions in the GQA dataset and 3 models: student model trained without augmentation, with augmentation and the auto-context teacher. Generated programs for existence questions about a single object and a single attribute like the one in the first column are almost always correct, even for the worse performing model, Llama-3.2-1B. In the second column, we see an example where the model trained without augmentation leaves out details mentioned in the question but the augmented model generates the correct program. All of the programs are incorrect in the last column but for different reasons. Both of the student models use the same program (apart from variable names) and make two mistakes. First the question asks about ‘not warm’ instead of ‘warm’ and the second is that the answer should be an object, not yes/no. The teacher program returns an object but still fails to recognize that the object should be ‘not warm’ even though the variable name includes ‘not warm’ in it. Questions involving negative properties are almost always missed by the teacher and student models.
Figure 4: 3 question/programs using no augmentation, augmentation and auto-context teacher. Simple comparison questions (left hand side) are almost always correct while questions with negations are almost always incorrect across the different methods.
Appendix B Training and Model Details
We used the following models for executing programs:
  1. 1.
    CLIP ViT-L/14 Radford et al. (2021)
  2. 2.
    InstructBLIP Flan-T5 XL Dai et al. (2023)
  3. 3.
    OWLv2 Base Patch 16 Ensemble Minderer et al. (2024)
Program generation settings for GPT can be found in Table 10. Template-based and direct training hyper-parameters can be found in Table 11. For CodeT5, the most important hyper-parameters were the learning rate and LoRA dropout rate. Training stopped when the training loss stopped decreasing.
All experiments were run on a single 40gb A40 or 40gb A100. Time measurements were measured on an A40.
Setting Value
Temperature 0
Top_p 1.0
Frequency Penalty 0.0
Presence Penalty 0.0
Max Output Tokens 256
Table 10: GPT-4o-mini generation settings
Hyper-parameters Value
LoRA target modules All linear layers
LoRA rank 8
LoRA alpha 16
LoRA bias None
LoRA dropout 0.05 (no augmentation), 0.1 (augmentation)
LR 2e-4
Batch Size 16
Max Output Tokens 256
Table 11: Training and evaluation settings for CodeT5
Appendix C Changes to ViperGPT API
The following are major modifications made to the ViperGPT API Surís et al. (2023).
Figure 5: Difference in program annotations when a VLM is used as a backup model for the question ‘Is the ground blue or brown?’ The highlighted portion is removed from all program annotations used.
  1. 1.
    Program annotations were modified not to use a vision-language model (VLM) when the program fails (see Figure 5 for an example). In the original ViperGPT API, examples in the API included a line to directly query a VLM if other parts of the program failed such as when no object is found. The performance using the original ViperGPT code decreases considerably when the VLM backup lines are removed from the API as shown in Table 12.
  2. 2.
    An object is always returned by the object detector.
  3. 3.
    Program annotations did not include parts of the program that were shared among all examples.
  4. 4.
    Several new functions were added.
    1. (a)
      Verify Relationship: Given two objects and a relation, return a boolean whether the objects satisfy that relationship.
    2. (b)
      Choose Relationship: Given two objects, return the relationship between the two.
    3. (c)
      Crop Position: Crop part of the image based on a position relative to an object
Use of VLM Backup GQA-Test Dev
ViperGPT with VLM Backup 47.3
ViperGPT without VLM Backup 26.0
Table 12: Change in GQA test-dev accuracy using original ViperGPT API when not using a VLM when the program fails
Appendix D Prompt
Instructions
For each question provided, generate a Python program that includes a return statement. Assume that image_patch = ImagePatch(image) is already defined. The final output of the program should always be a string.
ImagePatch
Attributes
1. **cropped_image**
Type: array
Description: An array representing the cropped image.
2. **left**
Type: int
Description: The left border of the crop’s bounding box.
3. **lower**
Type: int
Description: The bottom border of the crop’s bounding box.
4. **right**
Type: int
Description: The right border of the crop’s bounding box.
5. **upper**
Type: int
Description: The top border of the crop’s bounding box.
Methods
1. **find(object_name: str) -> List[ImagePatch]**
Description: Returns a list of image patches containing the specified object.
Notes: find should not be the last operation in a program.
Examples:
image_patch.find(’chair’)
image_patch.find(’table’)
2. **crop_position(direction: str, reference_patch: ImagePatch) -> ImagePatch**
Description: Returns a new image patch in the specified direction relative to the reference_patch. Directions can include ’left’, ’right’, ’above’, ’below’, ’on’, ’in front’, etc.
Notes: The result of crop_position should not be immediately indexed on the next line. The second argument is always the original image_patch.
Examples:
image_patch.crop_position(’left’, image_patch)
image_patch.crop_position(’above’, image_patch)
3. **verify_property(property_name: str) -> bool**
Description: Returns True if the object contains the specified property; otherwise, False.
Notes: Can only be called on an image patch.
Examples:
image_patch.verify_property(’red’)
image_patch.verify_property(’running’)
4. **classify(options: Union[str, List[str]]) -> str**
Description: Given a category (e.g., ’color’, ’material’, ’furniture’) or a list of options, returns the best option for the image patch.
Notes: The input should not be ’object’.
Examples:
image_patch.classify([’red’, ’blue’])
image_patch.classify(’color’)
5. **simple_query(question: str) -> str**
Description: Answers questions about the image, especially ambiguous ones (e.g., ’Who is riding?’).
Examples:
image_patch.simple_query(’Who is riding?’)
General Functions
1. **filter_img(image_patches: List[ImagePatch], criteria: str) -> List[ImagePatch]**
Description: Filters the list of image patches based on the given criteria. The criteria can be an action, attribute, or object.
Examples:
filter_img(image_patches, ’red’)
filter_img(image_patches, ’running’)
2. **choose_relationship(patch1: Union[ImagePatch, List[ImagePatch]], patch2: Union[ImagePatch, List[ImagePatch]], relationships: Union[List[str], str]) -> str**
Description: Chooses the relationship that best matches the two patches from the provided options.
Examples:
choose_relationship(image_patch1, image_patch2, [’on top of’, ’next to’])
choose_relationship(image_patch1, image_patch2, [’left’, ’right’])
3. **verify_relationship(patch1: Union[ImagePatch, List[ImagePatch]], patch2: Union[ImagePatch, List[ImagePatch]], relationship: str) -> str**
Description: Returns ’yes’ or ’no’ based on whether the specified relationship holds between the two patches.
Examples:
verify_relationship(image_patch1, image_patch2, ’on top of’)
verify_relationship(image_patch1, image_patch2, ’left’)
4. **exists(patches: Union[ImagePatch, List[ImagePatch]]) -> bool**
Description: Checks whether any of the provided image patches exist.
Notes: If used as the last operation, it should be followed by bool_to_yesno().
Examples:
exists(image_patches)
5. **bool_to_yesno(value: bool) -> str**
Description: Converts a boolean value to ’yes’ or ’no’. Used to convert outputs of verify_property and exists.
Examples:
bool_to_yesno(exists(image_patches))
Here are some examples of how to write programs:
{examples}
Additional Notes
- You may utilize standard Python functions within your programs.
- Do not include comments.
- Only return the program.
- Do not define the function.
- Functions never return None.
- The last line of each program should be answer =
Appendix E Variable Renamer
class VariableRenamer(ast.NodeTransformer):
def __init__(self, skip_vars=None):
self.counter = 1 # For general variables (var1, var2, ...)
self.temp_counter = 1 # For comprehension/loop variables (temp_var_1, ...)
self.name_map = {}
self.skip_vars = set(skip_vars) if skip_vars else set()
def _new_name(self):
name = f"var{self.counter}"
self.counter += 1
return name
def _new_temp_name(self):
[…]
self.temp_counter += 1
return name
def rename_target(self, target):
"""Rename normal assignment or loop targets, skipping those in skip_vars."""
if isinstance(target, ast.Name):
if target.id in self.skip_vars:
return target
if target.id not in self.name_map:
self.name_map[target.id] = self._new_name()
target.id = self.name_map[target.id]
elif isinstance(target, (ast.Tuple, ast.List)):
for elt in target.elts:
self.rename_target(elt)
return target
def visit_Name(self, node):
if isinstance(node.ctx, (ast.Store, ast.Load, ast.Del)):
if node.id in self.skip_vars:
return node
if node.id in self.name_map:
node.id = self.name_map[node.id]
return node
def visit_Assign(self, node):
node.value = self.visit(node.value)
node.targets = [self.rename_target(t) for t in node.targets]
return node
def rename_within(self, node, old_name, new_name):
"""Recursively replace occurrences of old_name with new_name within the node."""
class NameReplacer(ast.NodeTransformer):
def visit_Name(self, n):
if n.id == old_name:
n.id = new_name
return n
replacer = NameReplacer()
return replacer.visit(node)
def visit_For(self, node):
# Enhanced handling for For loops to propagate renaming within the loop body.
if isinstance(node.target, ast.Name) and node.target.id not in self.skip_vars:
old_name = node.target.id
new_temp = self._new_temp_name()
node.target.id = new_temp
# Visit and rename within ’iter’, ’body’, and ’orelse’
node.iter = self.visit(node.iter)
node.body = [self.rename_within(self.visit(n), old_name, new_temp) for n in node.body]
if node.orelse:
node.orelse = [self.rename_within(self.visit(n), old_name, new_temp) for n in node.orelse]
else:
node.target = self.rename_target(node.target)
node.iter = self.visit(node.iter)
node.body = [self.visit(n) for n in node.body]
if node.orelse:
node.orelse = [self.visit(n) for n in node.orelse]
return node
def visit_While(self, node):
node.test = self.visit(node.test)
node.body = [self.visit(n) for n in node.body]
if node.orelse:
node.orelse = [self.visit(n) for n in node.orelse]
return node
def visit_ListComp(self, node):
for gen in node.generators:
if isinstance(gen.target, ast.Name) and gen.target.id not in self.skip_vars:
old_name = gen.target.id
new_temp = self._new_temp_name()
gen.target.id = new_temp
node.elt = self.rename_within(node.elt, old_name, new_temp)
gen.ifs = [self.rename_within(if_clause, old_name, new_temp) for if_clause in gen.ifs]
for inner_gen in node.generators:
inner_gen.target = self.rename_within(inner_gen.target, old_name, new_temp)
else:
gen.target = self.rename_target(gen.target)
gen.iter = self.visit(gen.iter)
node.elt = self.visit(node.elt)
for gen in node.generators:
gen.ifs = [self.visit(if_clause) for if_clause in gen.ifs]
return node
def visit_GeneratorExp(self, node):
for gen in node.generators:
if isinstance(gen.target, ast.Name) and gen.target.id not in self.skip_vars:
old_name = gen.target.id
new_temp = self._new_temp_name()
gen.target.id = new_temp
node.elt = self.rename_within(node.elt, old_name, new_temp)
gen.ifs = [self.rename_within(if_clause, old_name, new_temp) for if_clause in gen.ifs]
for inner_gen in node.generators:
inner_gen.target = self.rename_within(inner_gen.target, old_name, new_temp)
else:
gen.target = self.rename_target(gen.target)
gen.iter = self.visit(gen.iter)
node.elt = self.visit(node.elt)
for gen in node.generators:
gen.ifs = [self.visit(if_clause) for if_clause in gen.ifs]
return node
def visit_With(self, node):
for item in node.items:
if item.optional_vars and isinstance(item.optional_vars, ast.Name) and item.optional_vars.id not in self.skip_vars:
item.optional_vars.id = self._new_temp_name()
elif item.optional_vars:
item.optional_vars = self.rename_target(item.optional_vars)
item.context_expr = self.visit(item.context_expr)
node.body = [self.visit(n) for n in node.body]
return node
# Additional visitor methods for other constructs can be added here.
def format_assignments(source_code: str) -> str:
Remove spaces around the equals sign in single-line assignment statements
without altering multi-line assignments.
This function ensures that:
- Single-line assignments have no spaces around ’=’.
- Multi-line assignments are left intact to preserve code correctness.
[…]
formatted_lines = []
[…]
# Track the balance of parentheses, brackets, and braces
paren_balance = 0
for line in lines:
stripped_line = line.strip()
# Update paren_balance
paren_balance += line.count(’(’) - line.count(’)’)
paren_balance += line.count(’[’) - line.count(’]’)
paren_balance += line.count(’{’) - line.count(’}’)
# If paren_balance > 0, we’re inside a multi-line expression
if paren_balance > 0:
formatted_lines.append(line)
continue
# Attempt to match an assignment statement
match = assignment_pattern.match(line)
if match:
indent, var, expr = match.groups()
# Remove spaces around ’=’ and reconstruct the line
formatted_line = f"{indent}{var}={expr}"
formatted_lines.append(formatted_line)
else:
# Non-assignment lines are added directly
formatted_lines.append(line)
# Join the lines back into a single string
[…]
def replace_variables(code: str, convert_to_source: bool = True) -> Union[str, ast.AST]:
skip_list = {"image_patch", "answer"} # Variables not to rename
tree = ast.parse(code)
renamer = VariableRenamer(skip_vars=skip_list)
new_tree = renamer.visit(tree)
ast.fix_missing_locations(new_tree)
new_source = ast.unparse(new_tree)
formatted_source = format_assignments(new_source)
return formatted_source
Given a generated program we call the function ‘replace_variables’ which uses and abstract-syntax tree to rename variables both in and outside different types of loops.
Appendix F Variable Replacement
Algorithm 1 Argument Replacement
  Extract arguments per function
  for func in functions do
     for arg in arguments do
        if arg in category then
           Random sample from category
        else
           Random sample from generic object list
        end if
     end for
  end for
The general algorithm for replacing an argument in a program can be seen in Algorithm 1. For each named function or method in the API, we extract the arguments. For GQA, if the argument is already in a pre-defined catgory, we randomly sample from that category. Otherwise we randomly sample an object. Some example categories and options can be seen in Table 13. The process for VQA is similar except there is no pre-defined list. Instead we mask out the argument in the question and generate a replacement using BERT or BART if the argument is a phrase. For full questions, we use word tokenization (default NLTK Tokenization) and POS-tagging with NLTK Bird et al. (2009) to determine where to place masks. We randomly sample from the top-50 results. Both the BERT and BART models are large uncased with 340 and 406 M parameters.
We perform the process above if an argument is selected for replacement during training. Each argument has a probability $p=0.5$ to be selected.
Table 13: Argument Categories and Options
Category Name Category Examples
Color red, blue, green, yellow, purple, black, white, orange, pink, brown, gray, indigo, cyan magenta, tan, silver
Activities running, walking, snowboarding, flying, splashing, tossing, riding, standing, hugging hanging, breaking, pulling, decorating, facing, preparing pouring, pointing, laughing
Relation picking up, in front of behind, above, below next to, near, far away close , following, on top, beside, walking on, attached, left, right, diagonal

### v4

@@
-Template-Based Visual Program Distillation
+Visual Program Distillation with Template-Based Augmentation
 Abstract
-For users with limited computational resources, visual programming or prompting large language models (LLMs) to generate executable code for visual tasks, like visual question answering (VQA), remains largely inaccessible.
-Even with techniques such as distillation, adapting visual programming to smaller models or specific datasets is still quite challenging due to high annotation costs.
-We propose a low-cost visual program distillation method that can be used for models with fewer than 1 billion parameters and requires no human-generated program annotations.
+Adapting visual programming or prompting large language models (LLMs) to generate executable code for visual tasks like visual question answering (VQA) for specialized tasks or domains remains challenging due to high annotation and inference costs.
+We propose a low-cost visual program distillation method that can be used for models with at most 1 billion parameters and requires no human-generated program annotations.
 We achieve this through synthetic data augmentation based on decoupling programs into higher-level skills, called templates, and their corresponding arguments.
-Experimental results show that, with a relatively small amount of question/answer data, small language models can generate high-quality visual programs with the added benefit of much faster inference.
-Template-Based Visual Program Distillation
+Experimental results show that, with a relatively small amount of question/answer data, small language models can generate high-quality specialized visual programs with the added benefit of much faster inference.
+Visual Program Distillation with Template-Based Augmentation
 Michal Shlapentokh-Rothman   Yu-Xiong Wang   Derek Hoiem University of Illinois at Urbana-Champaign {michal5,yxw,dhoiem}@illinois.edu
@@
 (2023).
-Despite success in zero-shot settings, visual programming remains largely inaccessible due to high program generation costs, which come from the need for expensive hardware to run LLMs or using expensive proprietary APIs, as well as long inference time and high data costs for both human-generated program annotations and question/answer annotations.
+Visual programming offers greater adaptability and customization for specialized applications such as a personal visual navigation assistant compared to a single vision-language model (VLM).
+In-context learning with proprietary or closed-source LLMs can generate correct visual programs for targeted applications but at the cost of long inference time and compute (see Figure 1).
+In addition, generating in-context examples (i.e.
+visual programs written by hand) requires a significant amount of human effort.
 Previous efforts Khan et al.
-(2024) around generating visual programs on smaller LLMs have made some progress in addressing these concerns, but training and inference costs remain substantial.
-Our goal is to create a visual programming system with two key characteristics:
+(2024) have made some progress in adapting smaller, open-source LLMs for dataset-specific visual programs but still suffer from high training and data costs.
+Figure 1: Accuracy vs.
+Throughput for Visual Program Generation on GQA Generalist LLMs (teacher models) offer high accuracy at the cost of low throughput and large model size (proportional to marker size).
+With our template-based augmentation method, specialized distilled student models achieve comparable performance on answer accuracy with a small percent of question/answer data ($\approx 0.1\%$) and no human program annotations.
+Our goal is to create a specialized visual programming system with two key characteristics:
   1. 1.
-    Small Program Generator Programs should be generated by models with $<1$ billion parameters, enabling fast inference and use on a wide variety of hardware.
+    Small Program Generator Programs should be generated by models with $\leq 1$ billion parameters, enabling fast inference and use on a wide variety of hardware.
   2. 2.
-    Limited Human Annotations Minimize the number of required annotations, especially human-generated programs.
-To achieve this, our approach builds on the key observation that many visual programs share a common high-level structure – a unique aspect of visual programming.
-For example, the programs ‘Count the red chairs’ and ‘Count the green apples’ follow the same pattern but differ in their inputs.
-Inspired by this insight and recent advances in LLM distillation Hsieh et al.
-(2023), we propose a low-cost visual program distillation method.
-Our approach requires no human-generated example programs and uses only a small fraction of question/answer pairs (at most 1000).
-A teacher model leverages auto-context generation, where each generated program that produces the correct answer is added to the set of in-context examples.
-The resulting annotated programs form a dataset used to train a small language model (SLM) to generate visual programs efficiently.
+    Limited Human Annotations Minimize the number of required annotations, especially human-generated programs, for ease of adaptation.
 Method Task Approach Model Size Train Size Domain
 VisRep Khan et al. (2024) Program Gen. Self-Training 7B 10% ($\approx$10K) Visual Prog.
-Distill Step-by-Step Hsieh et al. (2023) Text Gen. LLM Distillation 220M-770M 12.5% Q/A, Reasoning
+Distill Step-by-Step Hsieh et al. (2023) Text Gen. LLM Distillation 220M-1B 12.5% Q/A, Reasoning
 CodePlan Sun et al. (2024) Program Gen. LLM Distillation 770M 100% General Coding
-Template-Based Aug. (Ours) Program Gen. LLM Distillation 770M 0.1% ($\approx$1K) Visual Prog.
+Template-Based Aug. (Ours) Program Gen. LLM Distillation 500M-1B 0.1% ($\approx$1K) Visual Prog.
 Table 1: Comparison of approaches for program generation and LLM distillation.
 Our method enables LLM distillation for visual programming using only a small fraction of the training dataset and minimizing annotation costs.
-During training, program annotations are synthetically generated through template-based augmentation.
-This strategy decouples high-level skills or plans, called templates, from question-specific concepts or arguments.
-For example, the template for ‘Count the red chairs’ is ‘find(arg1), verify property(arg2), count’ and the corresponding arguments are ‘red’ (arg1) and ‘chair’ (arg2).
-New question/program pairs are created by replacing these arguments in the question and program.
-By replacing ‘red’ with ‘green’ and ‘chairs’ with ‘apples,’ we generate both a new question and a corresponding program.
-Compared to directly training on the program annotations, training with template-based augmentation offers three key benefits: 1) better performance when less data is available, 2) increase in student/teacher prediction agreement, and 3) higher-quality generated programs.
+Our key insight in achieving such capabilities lies in decoupling the skill or procedure from the question-specific concept.
+We call the higher-level skills templates and the concepts arguments.
+For example, the programs “Count the red chairs" and “Count the green bananas" have the same template,‘find(arg1), verify property(arg2), count’, but different arguments: ‘red’ (arg1) and ‘chairs’ (arg2) vs ‘green’ (arg1) and ‘bananas’ (arg2) respectively.
+This decomposition facilitates creating synthetic examples by replacing arguments in the question and program, e.g.
+‘Count the red apples’ with corresponding argument substitution in the program.
+Given a small training dataset, we can use this process, which we refer to as template-based augmentation to increase concept diversity.
+Combining template-based augmentation with recent advances in LLM distillation Hsieh et al.
+(2023), we propose a low-cost visual program distillation method.
+Our approach requires no human-generated example programs and uses only a small fraction of question/answer pairs (at most $0.1\%$ of the training dataset).
+A teacher model leverages auto-context generation, where each generated program that produces the correct answer is added to the set of in-context examples.
+The resulting annotated programs form a dataset of question/program pairs that are augmented via the template-based method described above and used to train a small language model (SLM) that efficiently generates visual programs.
 Our method uses a low-cost teacher model (e.g.
 GPT-Mini-4o) and costs only around \$1 per dataset.
-In addition, the distilled models achieve much faster inference speed than existing LLMs.
-To validate our approach, we evaluate on frequently used visual question answering (VQA) datasets and compare with using human-generated annotations as well as non-augmented distillation.
-Moreover, we address an overlooked evaluation aspect by performing several types of evaluation, including human program verification, due to the noisiness of commonly used metrics.
-All together, our results show that it is possible with few question/answer pairs to distill the visual program ability of LLMs into a much smaller model that does not require in-context examples.
+In addition, the specialized distilled models achieve much faster inference speed (up to $\approx 30.8$x) than the teacher models.
+To validate our approach, we evaluate on widely adopted visual question answering (VQA) datasets and compare with using human-generated annotations, few-shot prompting and non-augmented distillation.
+Most existing work evaluates only on whether the generated answer is correct.
+We find additional insights by also evaluating the correctness of the generated programs (as judged by a human) and student/teacher agreement, revealing substantial discrepancies among the different metrics.
+We summarize our contributions as follows:
+  • Introduce a template-based augmentation framework that distills large, generalist visual programming systems into small, specialized ones with minimal human effort.
+Template-based augmentation increases performance across all evaluated metrics.
+  • Empirically show auto-context generation achieves the same or better performance as human provided annotations when prompting a teacher model.
+  • Discuss and evaluate different metrics: answer accuracy, student/teacher agreement, and program accuracy for visual program evaluation.
+Surprisingly, we find the rate of error in answers is up to 3.8 times higher than that of the programs, suggesting that future visual programming work should focus on more effective APIs.
 2 Related Works
@@
 Visual programming has been applied to many different domains and applications including visual question answering, video question answering, text-to-image generation, and robotics (Cho et al., 2024; Lu et al., 2024; Min et al., 2024; Liang et al., 2023).
-Unlike previous visual programming works, we only use small models ($<1$ billion parameters) to generate programs.
-Visual programming is different from both LLM tool use and LLM code generation since it requires both basic reasoning and knowledge skills as well as basic code generation.
-One of the challenging parts of visual programming comes from the execution and evaluation due to underlying models.
+Unlike previous visual programming works, we only use small models ($\leq 1$ billion parameters) to generate programs.
+One of the challenges that distinguishes visual programming from LLM tool use is the dependence of execution and evaluation on the output of underlying models.
 As noted in other works Khan et al.
-(2024), incorrect programs can return (after being executed) the correct answer even if the program is incorrect.
+(2024), incorrect programs can return (after being executed) the correct answer even if the program is incorrect, which motivates our investigation into different evaluation metrics.
+A concurrent work, Visual Unit Testing (ViUniT) Panagopoulou et al.
+(2025), generates visual unit tests for both visual program selection and self-training.
+ViUniT also evaluates the unit test quality and program correctness and reports similar differences (36%) between program and answer accuracy.
 Tool-Based Finetuning
-Compared to prompt-based methods, there are relatively few methods focused on improving program generation in visual programming or in the general field of tool-based LLMs through finetuning.
+Compared to prompt-based methods, there are relatively few methods focused on improving program generation in visual programming for specific tasks or in the general field of tool-based LLMs through finetuning.
 One of the main challenges is the lack of program annotations for input/output pairs.
@@
 For each question/answer pair in a pre-existing training dataset, an LLM is prompted to generate a corresponding program.
-If the program decreases the training loss (of the same LLM) then the program annotation is added to a new dataset.
+If the program decreases the training loss (of the same LLM), the program annotation is added to a new dataset.
 Then the same LLM is finetuned on the generated dataset.
-A similar approach is used in Chain-of-Thought (CoT) finetuning (Zhu et al., 2023), program correction methods such as VDebugger (Wu et al., 2024) and visual programming based finetuning methods (Khan et al., 2024) such as VisRep.
-LoRA (Hu et al., 2021) is frequently used when finetuning LLMs on the generated data.
+A similar approach is used in Chain-of-Thought (CoT) finetuning (Zhu et al., 2023), program correction methods for specific datasets such as VDebugger (Wu et al., 2024) and specialized visual programming methods (Khan et al., 2024), such as VisRep.
+LoRA (Hu et al., 2022) is frequently used when finetuning LLMs on the generated data.
 In VisRep, Khan et al.
-(2024) use a self-training approach similar to toolformer for visual programming.
+(2024) use a self-training approach similar to toolformer to finetunue an LLM for dataset-specific visual programming.
 Self-annotated programs are kept if the executed program returns the correct answer.
 While there are some similarities between VisRep and our work, the underlying setting is different.
-VisRep focuses on improving program generation on an existing LLM (7B) through self-training, while the goal of our work is to distill visual program generation from an LLM to a small model (<1B).
+VisRep focuses on improving program generation on an existing LLM (7B) through self-training for specialized tasks, while the goal of our work is to distill visual program generation from an LLM to a small model ($\leq 1$B) for specialized tasks.
 Our work also does not require sampling across specific question types or correcting by hand different program annotations during training.
@@
 (2023).
-Then during inference, augmented generation (RAG) Lewis et al.
-(2020)can be performed to select the most relevant ones.
+Then during inference, retrieval augmented generation (RAG) Lewis et al.
+(2020b) can be performed to select the most relevant ones.
 We follow such an approach, referred to as auto-context generation by  Tao et al.
 (2024) to generate in-context examples for the teacher model.
+Figure 2: An overview of our augmentation method.
+Programs are first separated into templates and argument, new arguments are selected and plugged back into the question/program pair.
+Templates are created by renaming variables and removing question specific concepts.
+One single teacher generated question/program pair can turn into hundreds of new question/program pairs.
+3 Method
+Preliminaries
+Following the notation in ViperGPT Surís et al.
+(2023), the visual programming objective is to generate a program $z=\pi(q,p)$ with a program generator $\pi$, input query $q$ and prompt $p$ such that when executed with an execution engine $\phi$ and corresponding visual input $x$, $\phi(x,z)$ returns the correct answer.
+Prompt $p$ contains an API and dataset-specific in-context examples.
+Given a teacher model, $\pi_{t}$, and a smaller student model ($\leq 1$ billion parameters) $\pi_{s}$, our goal is to distill visual program generation from the teacher to the student for a specific dataset using a minimal number of answer annotations (question/answer pairs).
+We do not have any human provided program annotations (question/program/answer triplets), but we do know the API used for the teacher model.
+There are three main steps to our approach: teacher annotation, data augmentation and student training.
+3.1 Teacher Generated Program Annotation
+We use GPT-4o-Mini Open AI (2024) as our teacher model since it is relatively cheap and has strong performance.
+Unlike traditional knowledge distillation, LLM based distillation requires the teacher (an LLM) to be given the appropriate context through in-context examples before starting annotation.
+Common practice is to have a human generate the examples.
+We follow a process known as auto-context generation Tao et al.
+(2024) to automatically generate such examples given answer annotations only.
+Auto-context generation follows a simple process similar to VisRep Khan et al. (2024):
+  1. 1.
+    Teacher model predicts a program using API and in-context examples (if any).
+  2. 2.
+    Generated program is evaluated.
+  3. 3.
+    If the answer returned by a generated program matches the ground truth answer, then the program is immediately added to the set of in-context examples. Otherwise the program is discarded.
+For efficiency, the in-context examples are sampled based on similarity to the question (i.e.
+RAG) Lewis et al.
+(2020b) once there are more than $x$ examples.
+We compute the cosine similarity with finetuned MP-Net-v2 Reimers and Gurevych (2019) between an input and all in-context examples and select the $x=50$ highest scoring examples.
 Template image_patch = ImagePatch(image)
@@
 We want to answer similar questions the same way and easily generate synthetic data.
-3 Method
-Preliminaries
-Following the notation in ViperGPT Surís et al.
-(2023), the visual programming objective is to generate a program $z=\pi(q,p)$ with a program generator, $\pi$, input query $q$ and prompt $p$ such that when executed with execution engine $\phi$ and corresponding visual input $x$, $\phi(x,z)$ returns the correct answer.
-Prompt $p$ contains an API and dataset-specific in-context examples.
-Figure 1: An overview of our augmentation method.
-Programs are first separated into templates and argument, new arguments are selected and plugged back into the question/program pair.
-Templates are created by renaming variables and removing question specific concepts.
-One single teacher generated question/program pair can turn into hundreds of new question/program pairs.
-Given a teacher model, $\pi_{t}$, and a smaller student model ($<1$ billion parameters), $\pi_{s}$, our goal is to distill visual program generation from the teacher to the student for a specific dataset using a minimal number of answer annotations (question/answer pairs).
-We do not have any human provided annotations (question/program/answer triplets) but we do know the API used for the teacher model.
-There are three main steps to our approach: teacher annotation, data augmentation and student training.
-3.1 Teacher Generated Program Annotation
-We use GPT-4o-Mini AI (2024) as our teacher model.
-Unlike traditional knowledge distillation, LLM based distillation requires the teacher (an LLM) to be given the appropriate context through in-context examples before starting annotation.
-Common practice is to have a human generate the examples.
-We follow a process known as Auto-Context Generation to automatically generate such examples given answer annotations only.
-Figure 2: An example of our data augmentation approach.
-Both the new and old question have the same template so the template matcher output should predict the same template for both.
-The arguments for the new and old programs are different.
-But, in the arguments, (dog, sofa, brown) should be replaced with (bear, desk, green).
-Auto-context generation follows a simple process similar to VisRep Khan et al. (2024):
-  1. 1.
-    Teacher model predicts a program using API and in-context examples (if any).
-  2. 2.
-    Generated program is evaluated.
-  3. 3.
-    If the answer returned by a generated program matches the ground truth answer, then the program is immediately added to the set of in-context examples. Otherwise the program is discarded.
-For efficiency, the in-context examples are sampled based on similarity to the question (i.e.
-RAG) Lewis et al.
-(2020) once there are more than 50.
-We compute the cosine similarity with finetuned MP-Net-v2 Reimers and Gurevych (2019) between an input and all in-context examples and select the 50 highest scoring examples.
 3.2 Data Augmentation
@@
 The general structure of each program is the same with the only difference coming from the inputs to the functions.
-If we answer one of these questions correctly and know that the remaining questions have the same structure, then all of the remaining questions should have that same structure or should be consistent.
+If we answer one of these questions correctly and know that the remaining questions have the same structure, then all the remaining questions should also have that same structure or be consistent.
 An overview of our data augmentation method can be seen in Figure 2.
-We refer to this as ‘template-based augmentation.’
+We refer to this as "template-based augmentation."
+Example Template
+Is the blue car the same shape as the chair? Is leather jacket made of the same material as the shirt? image_patch = ImagePatch(image)
+var1 = image_patch.find(<arg_0>)
+var2 = filter_img(<arg_1>)
+var3 = var2.classify(<arg_2>)
+var4 = image_patch.find(<arg_3>)
+var5 = var4.classify(<arg_4>)
+answer =bool_to_yesno(var3 ==== var5)
+What type of food is near the person? What is the vehicle next to the animal? image_patch = ImagePatch(image)
+var1 = image_patch.find(<arg_0>)
+var2 = var1.crop_position(<arg_1>)
+var3 = var2.find(<arg_2>)
+answer == var3.classify(<arg_3>)
+Is the car to the left or right of the tree? Is pot above or below the pan? image_patch = ImagePatch(image)
+var1 = image_patch.find(<arg_0>)
+var2 = image_patch.find(<arg_1>)
+answer == choose_relationship(var1,var2,<arg_2>)
+Table 3: Some examples of questions and corresponding templates. Multi-colored words correspond to multiple arguments.
 Templates
-We define a template as a specific ordering of functions, where a function is an API call to a visual model or python operation. Templates are argument independent. For example, if the program is
+We define a template as a specific ordering of functions, where a function is an API call to a visual model or python operation.
+Templates encode program structure independently of argument values.
+For example, if the program is
 image_patch = ImagePatch(image)
@@
 (2024).
-The main advantage to templates is that they can be extracted directly from the program.
+The main advantage of templates is that they can be extracted directly from the program.
 There is no need to generate extra output or perform multiple forward passes for a single question.
+Figure 3: An example of our data augmentation approach.
+Both the new and old question have the same template, so the template matcher output should predict the same template for both.
+The arguments for the new and old programs are different.
+But, in the arguments, (dog, sofa, brown) should be replaced with (bear, desk, green).
 Template Extraction
@@
 The variable renaming and extraction can be done in seconds with abstract syntax trees and regular expression matching.
-The code can be found in Appendix E.
+Note that while the template extraction algorithm is deterministic, the actual templates are determined solely by the annotations from the teacher model.
+The code can be found in Appendix F.
 Augmentation
@@
 As can be seen in Table 2, for many questions and programs, the arguments appear directly in the question.
-Consider the example in Figure 2: “Are both the dog and the sofa brown?" The arguments are dog, sofa, and brown.
+Consider the example in Figure 3: "Are both the dog and the sofa brown?" The arguments are dog, sofa, and brown.
 Once we find similar words for each, we can simply replace them in the sentence.
-Since we already know the template and the arguments, we also have the program.
-For GQA (Hudson and Manning, 2019), the possible word replacements are from ViperGPT Surís et al.
-(2023) and for arguments that are either not in the question or are equal to the entire question, we use GPT to generate similar questions.
+At this point, we also have the program, since the template and arguments are known.
+There are two different methods for generating word replacements.
+For the GQA (Hudson and Manning, 2019) dataset, the possible word replacements come from ViperGPT Surís et al.
+(2023) except for two special circumstances.
+The first is when arguments are not in the question.
+For example, the question "What is that made from?" could have a program where the arguments are "object" (with the find function) and "material" (with the classify function).
+The second is when the program consists of only calling a VLM and the input is the entire question.
+For the special cases, we use GPT Open AI (2024) to generate similar questions.
 For VQAv2 Goyal et al.
 (2017), we use BERT Devlin et al.
-(2019) when an argument is a single word and BART Lewis (2019) to replace phrases.
-If possible, we replace arguments with arguments of the same type e.g.
+(2019) when an argument is a single word and BART Lewis et al.
+(2020a) to replace phrases.
+If possible, we replace arguments with arguments of the same type, e.g.
 an attribute is replaced by an attribute.
-Example Template
-Is the blue car the same shape as the chair? Is leather jacket made of the same material as the shirt? image_patch = ImagePatch(image)
-var1 = image_patch.find(<arg_0>)
-var2 = filter_img(<arg_1>)
-var3 = var2.classify(<arg_2>)
-var4 = image_patch.find(<arg_3>)
-var5 = var4.classify(<arg_4>)
-answer =bool_to_yesno(var3 ==absent$=$ = var5)
-What type of food is near the person? What is the vehicle next to the animal? image_patch = ImagePatch(image)
-var1 = image_patch.find(<arg_0>)
-var2 = var1.crop_position(<arg_1>)
-var3 = var2.find(<arg_2>)
-answer $=$ var3.classify(<arg_3>)
-Is the car to the left or right of the tree? Is pot above or below the pan? image_patch = ImagePatch(image)
-var1 = image_patch.find(<arg_0>)
-var2 = image_patch.find(<arg_1>)
-answer $=$ choose_relationship(var1,var2,<arg_2>)
-Table 3: Some examples of questions and corresponding templates. Multi-colored words correspond to multiple arguments.
 During training, each argument in each question has a 50% chance of being replaced.
@@
 fewer than 10).
-Some examples of categories and more details on word replacement can be found in Section F.
+Some examples of categories and more details on word replacement can be found in Appendix G.
 Student Training
 Given the augmented dataset, we perform LoRA-based finetuning on the student model.
-Since the augmentation method produces both questions and corresponding programs, we use next-token prediction loss for training.
+The input to the student model is the question and the output is the visual program.
+Since the augmentation method produces both questions and corresponding programs, we use a next-token prediction loss for training.
 4 Experiments
 Experimental Setup
-For all experiments, the teacher model is GPT-4o-mini, the student model is CodeT5 Wang et al.
-(2023) and is trained with LoRA Hu et al.
+For all experiments, the teacher model is GPT-4o-mini.
+Several student models are trained with LoRA Hu et al.
+(2022): Qwen2.5-Coder-0.5B Hui et al.
+(2024), CodeT5 Wang et al.
+(2023) and Llama-3.2-1B Dubey et al.
 (2024).
 We use the GQA dataset Hudson and Manning (2019) and VQAv2 Goyal et al.
-(2017) datsets for our experiments.
-We evaluate on the full GQA test-dev split and randomly sample 10000 questions from the VQAv2 validation split.
+(2017) datasets for our experiments.
+All student models are trained and evaluated on a single dataset unless indicated by the word ‘joint’.
+We evaluate on the full GQA test-dev split and randomly sample 10,000 questions from the VQAv2 validation split.
 GQA is evaluated using exact match and VQAv2 is evaluated based on annotator/answer agreement as in the original benchmark.
-We use a slightly modified API from the original ViperGPT paper  Surís et al.
-(2023).
-The main differences are some additional functions (to reduce program length) and removal of the use of a VLM if earlier parts of a program fail.
-Please see Appendix Section C for more details.
-For visual models, we use InstructBLIP (Flan-T5 XL) Dai et al.
-(2023) for general visual queries, Owl-ViT2 Minderer et al.
-(2024) for detection and CLIP Radford et al.
-(2021) for classification.
-Auto-Context Generation
-Before training the student model, we adapt the teacher to the VQA domain.
-In Table 4, we investigate the effectiveness of using human-generated programs and programs generated by the teacher model validated by answer correctness.
-As our aim is to enable visual programming with low cost and effort, we compare using a small number (25) of human-generated programs to generating 1000 programs (yielding hundreds of validated programs).
-Dataset Human Generated Auto-Context Generated Performance
+Experiments on additional datasets can be found in Appendix A.
+Dataset Human Generated LLM Generated Performance
 GQA 0 0 35.1
@@
 25 0 54.2
+0 286 (out of 500) 60.3
 25 283 (out of 500) 57.8
-0 286 (out of 500) 60.3
-Table 4: Teacher performance compared across varying numbers of human-generated and auto-context-generated in-context examples.
+Table 4: Teacher performance compared across varying numbers of human-generated and LLM generated in-context examples (i.e.
+auto-context generation).
+Auto-Context Generation achieves the same or better performance than human-generated program annotations.
+Error Example Explanation
+Not executable Q: Is the chair left or right?
+P: side = choose_relationship (chair, img, [‘left’,‘right’])
+choose_relationship requires a list as input
+API Violation Q: What color is the running dog on the left?
+P: crop_position (‘running left’,img)
+“running left” isn’t a valid direction/preposition.
+Contradicts Question Q: What color is the car above the road?
+P: below_road = crop_position(‘below’, road)
+The question states that the car is above the road, not below.
+Does Not Answer Question Q: Are there two tables?
+P: num = count(image.find(‘table’))
+The question asks if there are two tables, not how many tables there are.
+Does Not Include All Question Information Q: Is the blue toy small?
+P: toy.verify_property(‘small’)
+The question specifically asks about the blue toy.
+Table 5: Common errors encountered during qualitative evaluation.
+We use a slightly modified API from the original ViperGPT paper  Surís et al.
+(2023).
+The main differences are some additional functions (to reduce program length) and removal of the use of a VLM if earlier parts of a program fail.
+Please see Appendix D for more details.
+For visual models, we use InstructBLIP (Flan-T5 XL) Dai et al.
+(2023) for general visual queries, Owl-ViT2 Minderer et al.
+(2024) for detection and CLIP Radford et al.
+(2021) for classification.
+Method GQA VQA
+Answer Acc Prog. Acc Stu/Tea Agree. Answer Acc Prog. Acc Stu/Tea Agree.
+Few-shot prompting
+Llama 3.2-1B 19.0 22 23.3 40
+Qwen Coder 2.5-0.5B 18.1 14 25.0 16
+Teacher
+GPT-4o-Mini 43.1 85 60.3 93
+Distilled Students w/Aug.
+Qwen Coder 2.5-0.5B (single) 42.5 71 78.8 60.8 73 74.2
+Qwen Coder 2.5-0.5B (joint) 43.1 78 78.2 60.3 79 70.5
+Code-T5-770M (single) 41.9 68 64.5 61.1 80 76.1
+Llama 3.2-1B (single) 43.1 73 81.2 60.3 82 73.6
+Table 6: Comparison of few-shot prompting and student models trained w/ augmented distillation across different metrics.
+Few-shot open, source models have poor performance but template-augmented student distillation has comparable performance to the teacher model, GPT-4o-Mini.
+Auto-Context Generation
+Before training the student model, we adapt the teacher to the VQA domain.
+We investigate the effectiveness of using human-generated programs and programs generated by the teacher model validated by answer correctness.
+As our aim is to enable visual programming with low cost and effort, we compare using a small number (25) of human-generated programs to generating 1,000 programs (yielding hundreds of validated programs).
 As shown in Table 4, human-generated programs provide significant improvement vs.
@@
 This finding supports using self-generation to improve the teacher model, as question-answer pairs require much less expertise and time to provide.
-Data Usage GQA VQAv2
-Aug. No Aug. Aug. No Aug
-0.1% 41.9 (+0.8) 41.1 61.1 (+0.3) 60.8
-0.05% 39.7 (+4.1) 35.6 59.9 (+9.1) 50.8
-0.02% 34.2 (+4.2) 30.0 49.0 (+7.8) 41.2
-Table 5: Answer accuracy with and without augmentation across different dataset sizes: 1000 (0.1%), 500 (0.05%), and 250 (0.02%) for GQA and 500 (0.1%), 250 (0.05%) and 125 (0.02%) for VQAv2.
-Performance gain from augmentation is higher when using 0.05% and 0.02% of the dataset compared to 0.1%.
-Augmentation vs Non-Augmentation
-When training the student model, we use templates and variable substitution to create a more diverse set of programs for training.
-We evaluate the effect of template-based data augmentation in two ways.
-In Table 5, we analyze answer accuracy as a function of varying dataset sizes.
-The results indicate that data-augmentation increases the answer accuracy of the student model, with the largest benefits when few generated programs are available for training.
-To better evaluate the effectiveness of distillation, we also evaluate whether the teacher and student provide the same answer.
-In Table 6, we see that augmentation leads to much higher rates in student-teacher agreement, even when the overall improvement in accuracy is small (see Table 5).
-GQA VQA
-With Augmentation 64.5 76.1
-No Augmentation 59.8 72.0
-Table 6: Student/Teacher Prediction Agreement. Data augmentation has a much larger effect on student/teacher prediction compared to answer accuracy.
-Cost and Efficiency
-One of our objectives is keeping annotation costs low.
-As shown in Table 7, the total annotation cost for program auto-generation on both datasets is less than a dollar.
-GQA costs a bit more because of the larger size and during augmentation, GPT-4o-Mini is called when arguments do not appear in the question.
-By using a small model, inference speed also greatly increases.
-GQA VQAv2
-Annotation Cost \$0.69 \$0.26
-CodeT5 Inference Time (q/s) 35 32.7
-Llama-3.2-1B Inference Time (q/s) 0.32 0.33
-GPT-4o Mini Inference Time (q/s) 1.3 1.3
-Table 7: Annotation cost and inference time for several different models. Our method has low cost and high inference speed with CodeT5.
-Comparison with Related Work and Program Evaluation
-We compare our approach with the most closely related setting: prompt-based visual program generation.
-Other works, such as VisRep Khan et al.
-(2024), have different objectives and use carefully curated and larger datasets, and thus are not included in our comparisons.
-Furthermore, VisRep evaluates on selected subsets of various datasets, and the specific subsets, models, and code have not been made publicly available for comparison.
-Method GQA VQA
-Answer Acc Program Acc Answer Acc Program Acc
-GPT-4o-Mini (w/Human Examples) 38.5 78 54.2 92
-Llama 3.2 1B (w/Human Examples) 19.0 22 23.3 40
-Code-T5 Distilled w/o Augmentation (Ours) 41.1 48 60.8 79
-Code-T5 Distilled w/Augmentation (Ours) 41.9 68 61.1 80
-GPT-4o-Mini (Auto-context) (Ours) 43.1 85 60.3 93
-Table 8: Answer and program accuracy of different methods on GQA and VQAv2.
-With augmentation, distilled programs can achieve similar or better answer accuracy as prompting methods and improve program accuracy compared to not using augmentation.
-The widely followed ViperGPT Surís et al.
-(2023) approach is to use a prompt containing an API and a set of human generated in-context examples.
-We use two models to evaluate prompting GPT-4o-Mini and Llama-3.2-1B Dubey et al.
-(2024), a distilled version of Llama 3.
-25 human generated in-context examples are used per dataset for each model.
-In Table 8, we refer to these baselines as GPT-4o-Mini (w/human examples) and Llama-3.2-1B (w/human examples).
-From our method, we evaluate on two distilled CodeT5 models with and without augmentation.
-The distilled models are trained with 0.1% of the data (474 question/answer pairs for GQA and 286 question/answer pairs for VQAv2).
-We also include the auto-context teacher model, notated as GPT-4o-Mini (Auto-context).
-Table 9: Common errors and examples used during qualitative evaluation.
-Error Question Program Explanation
-Not executable Is the chair on the left or right side? chair = image_patch.find(‘chair’) side = choose_relationship(chair, image_patch, ‘left or right’) choose_relationship requires a list as input
-API Violation What color is the running dog on the left?
-left = crop_position(‘running left’, image_patch) dog = left.find(‘dog’) color = dog.classify(‘color’) ‘Running left’ is not a direction or preposition
-Contradicts Question What color is the car above the road?
-road = image_patch.find(‘road’) below_road = crop_position(‘below’, road) car = below_road.find(‘car’) car_color = car.classify(‘color’) The question states that the car is above the road, not below.
-Does Not Answer Question Are there two tables? tables = image_patch.find(‘table’) num_tables = count(tables) The question asks if there are two tables, not how many tables there are.
-Does not include all question information Is the blue toy small?
-toy = image_patch.find(‘toy’) is_small = toy.verify_property(‘small’) answer = is_small The question specifically asks about the blue toy.
-We evaluate using two metrics: answer accuracy and program accuracy.
+Evaluation Metrics
+All approaches are evaluated with two metrics: answer accuracy and program accuracy.
+Student models are also evaluated on student/teacher answer agreement or if student and teacher programs return the same answer, regardless of correctness.
 Program accuracy involves a human (the authors) manually evaluating each program (without execution) for correctness.
-We randomly sample 100 questions from each dataset for evaluation.
+We randomly sample 100 questions from each dataset for program correctness evaluation.
 The same 100 questions are used to evaluate each method.
+To ensure the fairness of our program evaluation, we enlist a second annotator to evaluate program correctness on 50 examples: 25 correct and 25 incorrect.
+The second annotator agreed with the original program evaluation on 47 out of 50 of the examples.
+Program annotations have been released so others can contribute.
+[^1] https://github.com/michalsr/vp_distill_templates/tree/main/program_correctness_annotations
 There are multiple ways to determine if a program is correct. We generally assume a program is correct unless it violates one of these criteria:
-  • Not Executable: The program must be executable and return the correct data type (a string for VQA datasets).
-  • API Violation: Visual programming APIs are designed to follow basic visual knowledge and reasoning.
+- Not Executable The program must be executable and return the correct data type (a string for VQA datasets).
+- API Violation Visual programming APIs are designed to follow basic visual knowledge and reasoning.
 For example, the ‘find’ function is used with nouns while ‘verify_property’ is generally used for attributes.
@@
 ‘crop_position(running)’) are considered incorrect.
-  • Contradicts Question: Programs that assume a statement that directly conflicts with a statement in the question.
+- Contradicts Question Programs that assume a statement that directly conflicts with a statement in the question.
 For example, assuming an object is on the left, when the question states it is on the right.
@@
 Some examples can be seen in Table 5.
-We can draw several observations from the results in Table 8.
-From the answer accuracy results, we see that with a small amount of data and a strong enough teacher, student models can achieve similar performance to the auto-context trained teacher.
-Auto-context generation also improves the teacher performance.
-The performance from prompting Llama is particularly low, consistent with previous findings showing that it is difficult to transfer prompts across LLMs Sclar et al.
-(2024) of similar size, let alone much smaller ones.
-The program accuracy results are bit more surprising.
-For all of the methods, program accuracy is higher than answer accuracy especially for GPT-4o-Mini (Auto-context) which has 85% vs 43% and 93% vs 60% accuracy for GQA and VQAv2 (respectively).
-Data augmentation has a larger effect on program accuracy than question accuracy but a substantial gap still remains between the student and teacher models, indicating more room for improvement.
-Across the two datasets, we see a much higher overall program accuracy for VQAv2 than for GQA.
-Both datasets have fairly easy questions but VQAv2 has many that are correctly answered by either calling a vision-language model (the simple_query) function or simple counting, which involves two functions (find and count).
-Taken together, our evaluation suggests that the quality of the visual programs is much better than what answer accuracy alone would imply.
-We include additional analysis of program errors in Appendix A.
+Few-Shot Comparison
+First, we compare our distilled student models with the most closely related setting: prompt-based visual program generation.
+Other works, such as VisRep Khan et al.
+(2024), have different objectives and use carefully curated and larger datasets, and thus are not included in our comparisons.
+Furthermore, VisRep evaluates on selected subsets of various datasets, and the specific subsets, models, and code have not been made publicly available for comparison.
+Two models, Llama-3.2-1B and Qwen2.5-Coder-0.5B, are evaluated using both few-shot prompting and template-based distillation.
+CodeT5 is also evaluated with template-based distillation.
+There are two versions of distilled Qwen2.5-Coder-0.5B: one version is trained (and evaluated) on GQA and VQA separately (denoted by ‘single’) and the other is trained on both datasets (denoted by ‘joint’).
+The same 25 human generated in-context examples (incorporated into the ViperGPT Surís et al.
+(2023) API) used in the Auto-Context Generation experiments above are used per dataset for each few-shot model.
+The distilled models are trained with 0.1% of the data (474 question/answer pairs for GQA and 286 question/answer pairs for VQAv2).
+All program annotations come from the teacher model.
+Method GQA VQA
+Answer Acc Prog. Acc Stu/Tea Agree. Answer Acc Prog. Acc Stu/Tea Agree.
+Qwen Coder 2.5-0.5B w/o Aug (single) 41.8 69 73.0 60.2 75 72.2
+Qwen Coder 2.5-0.5B w/ Aug (single) (+0.7)42.5 (+2)71 (+5.8)78.8 (+0.6)60.8 (-2)73 (+2.0)74.2
+Qwen Coder 2.5-0.5B w/o Aug (joint) 41.7 69 73.5 58.8 72 67.6
+Qwen Coder 2.5-0.5B w/ Aug (joint) (+1.4) 43.1 (+9)78 (+4.7)78.2 (+1.5)60.3 (+6)78 (+2.9)70.5
+Code-T5-770M w/o Aug (single) 41.1 48 59.8 60.8 79 72.0
+Code-T5-770M w/ Aug (single) (+0.8)41.9 (+20)68 (+4.7)64.5 (+0.3)61.1 (+1)80 (+4.1)76.1
+Llama 3.2-1B w/o Aug (single) 40.0 61 70.1 60.2 79 72.6
+Llama 3.2-1B w/ Aug (single) (+3.1)43.1 (+12)73 (+11.1)81.2 (+0.1)60.3 (+3)82 (+1)73.6
+Table 7: Effect of augmentation.
+On average, template-based augmentation improves performance on nearly all models, particularly on student/teacher agreement and program accuracy.
+Bold results are statistically significant.
+Few-Shot Results We can draw several observations from the results in Table 6.
+Comparing different methods, we see that few-shot prompting on small open-source models results in extremely low performance but with a small amount of data and a strong enough teacher, the same models achieve similar performance to the auto-context teacher, especially on answer accuracy.
+Program accuracy improves significantly from distillation but a gap still remains indicating room for improvement.
+The results across different metrics are a bit more surprising.
+On all of the student models, the average difference between student/teacher answer agreement and answer accuracy is 33% for GQA and 13% for VQAv2, indicating that answer accuracy underestimates distillation performance.
+The noisiness of answer accuracy is illustrated even more by program accuracy performance.
+The program accuracy results are the most surprising.
+For all of the methods, program accuracy is higher than answer accuracy especially for the teacher model GPT-4o-Mini which has 85% vs.
+43% and 93% vs.
+60% program vs answer accuracy for GQA and VQAv2 (respectively).
+The significant difference among the metrics indicates that most errors in visual programming systems with proprietary LLMs are not from the programs but from the API and visual models.
+An additional analysis of program errors is in Appendix B.
+Data Augmentation
+Next, we ablate the effect of data augmentation on distillation.
+For each model, we train with and without distillation on both datasets and measure performance across the three metrics.
+From the results in Table 7, we see that data augmentation improves performance across all metrics for nearly all models and datasets.
+Augmentation has more of an effect on GQA compared to VQAv2.
+Both datasets have fairly easy questions, but VQAv2 has many questions that are correctly answered by either calling a vision-language model (the simple_query) function or simple counting, which involves two functions (find and count).
+The relative increase is more notable for program accuracy (average 6.4) and student/teacher agreement (average 4.5) compared to answer accuracy (average 1.1), again indicating that answer accuracy does not fully capture model behavior.
+Cost and Efficiency
+One of our objectives is low annotation costs and fast inference time.
+As shown in Table 8, the total annotation cost for program auto-generation on both datasets is less than a dollar.
+GQA costs a bit more because of the larger size and during augmentation, GPT-4o-Mini is called when arguments do not appear in the question.
+By using a small model, inference speed also greatly increases.
+Metric GQA VQAv2
+Teacher
+Program Generation Cost (Total) \$0.69 \$0.26
+GPT-4o-Mini Inference Time (q/s) 1.3 1.3
+Distilled Students
+Qwen2.5-Coder Inference Time (q/s) 39.2 40.4
+CodeT5 Inference Time (q/s) 35.0 32.7
+Llama-3.2-1B Inference Time (q/s) 16.1 18.6
+Table 8: Annotation cost and inference time for student and teacher models. Student inference time is much faster than teacher inference time.
 5 Conclusion
-Our experiments show with template-based visual program distillation, only a small number of answer annotations are needed to achieve similar performance to prompt-based methods.
+Our experiments demonstrate models trained with template-based visual program distillation can become specialized and efficient visual program generators at a small cost.
 Auto-context generation removes the burden of human generated program annotations while still retaining the same performance.
-Template-based augmentation improves both student/teacher prediction agreement and program accuracy compared to non-augmented approaches.
-The distilled student models achieve similar answer accuracy as the teacher at a fraction of the time and cost.
-In conclusion, we show that the visual program ability of LLMs can be distilled into much smaller models with only a few question/answer pairs and no human-generated programs.
-For less than \$6, (\$5 for student training on the cloud and \$1 for annotation), a 770-M coding model can become a visual program generator.
-We anticipate that use of template-based visual program distillation will enable users and researchers to iterate more quickly on various visual programming systems and broaden their applications.
-6 Limitations
-There are several limitations to our work:
-Indirect Evaluation Answer accuracy is an indirect evaluation of program correctness.
-As indicated by our program accuracy results shown in Section 4, answer-based evaluation is only a loose indicator of visual program correctness.
-Promising areas for future work include automatic program accuracy evaluation and closing the gap between answer accuracy and program accuracy.
-Reliance on teacher model and API Our method relies on the quality of the teacher model, which in turn is dependent on the quality of the API and general prompt.
+The results also show how commonly used metrics for visual programming, do not fully capture the performance.
+Human program verification reveals that on the best models, programs are likely not the source of errors and that future work should focus on the API and visual models, not program generation.
+For less than \$6, (\$5 for student training on the cloud and \$1 for annotation), a 500-M coding model can become a visual program generator.
+We anticipate that the use of template-based visual program distillation will enable users and researchers to iterate more quickly on various visual programming systems and broaden their use for targeted applications.
+6 Limitations and Future Work
+There are several limitations and areas for future work:
+Specialization.
+Our method distills a specialized VQA model that is not intended to provide the same breadth of capability as the original LLM.
+Our specialized models, however, are much faster and cheaper for inference and easy to produce, making them suitable for targeted application.
+Effort Required for Program Accuracy.
+While program accuracy is an important metric, it also requires a significant amount of human effort, making it difficult to evaluate on a large scale.
+A promising area for future work includes automatic evaluation of program accuracy.
+Reliance on teacher model and API.
+Our method relies on the quality of the teacher model, which in turn is dependent on the quality of the API and general prompt.
 A future challenge is how to learn from a weaker teacher and/or unreliable API.
-Teacher Data Efficiency Following prior works, if a program was incorrect, it was not used.
+Teacher Data Efficiency.
+Following prior works, if the answer returned by a program was incorrect, it was not used.
 However, LLMs have the ability to self-correct given the appropriate feedback.
@@
 (2024).
-Program Execution Time Our method decreases program generation time, but not program execution time.
-With a single GPU, program execution time on average was 3.37 s and varies quite widely.
-The long execution time remains a bottleneck in different applications of visual programming.
-Limited Program Complexity Existing VQA datasets are fairly simple and most work on visual programming is limited to tasks where the programs can be generated in a single step by an LLM.
+Program Execution Time.
+We evaluate the time to generate programs, but the time to execute with API calls to several visual models, can be much greater (3.4s / query in our implementation with high variance) and requires significant engineering effort to make efficient.
+Limited Program Complexity.
+Existing VQA datasets are relatively simple and most work on visual programming is limited to tasks where the programs can be generated in a single step by an LLM.
 Most real world applications are multi-step and would require more complex reasoning and knowledge skills than in the evaluation datasets.
+7 Societal Implications
+Distilled student models inherit existing biases from LLMs so care is required when deploying the models in production.
+There are also privacy risks, so safeguards should be taken to prevent data leakage.
 Acknowledgement
@@
 (2023), which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
-Appendix A Program Analysis
+Appendix A Additional Results
+Additional experimental results on two vision-centric datasets, the 2D-split of CV Bench Tong et al.
+(2024) and the VizWiz Gurari et al.
+(2018) validation set, can be found in Table 9.
+We evaluate an augmented student model (Qwen2.5-Coder-0.5B) and the teacher model.
+For CV Bench Tong et al.
+(2024), the student model is trained on both GQA and VQAv2 (referred to as ‘joint’ in Section 4).
+2000 examples ( 0.1%) of the VizWiz training set are annotated by the teacher model and used to train the student.
+Like the experiments on VQAv2, we use BERT Devlin et al.
+(2019) and BART Lewis et al.
+(2020a), for augmentation.
+The results show the efficacy of our method on datasets related to the training set as well as real-world tasks.
+Method CV Bench VizWiz
+Answer Acc Stu/Tea Agree. Answer Acc Stu/Tea Agree.
+GPT-4o-Mini (Teacher) 56.0 N/A 34.3 N/A
+Qwen Coder 2.5-0.5B (w/ Aug) 54.5 93.0 34.9 71.6
+Table 9: Additional evaluation of augmented distillation on the 2D split of CV Bench and the validation split of the real-world dataset VizWiz.
+The distilled augmented model has similar performance as the teacher model.
+N-Gram Entropy
+In our experiments, we primarily use task-level metrics to validate our method and assess augmented data quality.
+To directly quantify data diversity, we also measure N-Gram Entropy Zhang et al.
+(2018) on 20,000 samples from both our augmented and non-augmented GQA training sets.
+The augmented data has an entropy of 7.10 compared to 6.18 for the original set, confirming that our augmentation procedure meaningfully increases linguistic variety.
+Probability of Augmentation
+In Section 3.2, we state that during augmentation, each argument has a 50% probability of being replaced.
+We evaluate using different percentages on a small validation set of GQA and found 50% had the highest performance, 43.2%, compared to 42.5% for 90% replacement rate and 42.4% for a 10% replacement rate.
+Model Size
+The focus of our work is to design a method for a small amount of training data as well small models that can be run on consumer GPUs, which led us to focus on models with no more than 1B parameters.
+However, there are models that are slightly larger, such as 1.5B which leads to the question, whether there are any performance gains from using such models.
+On a subset of GQA, we evaluate different sizes of Qwen2.5-Coder on 0.5B, 1.5B and 3B.
+The results, shown in Table 10, show our augmentation method is actually more effective on smaller models.
+Scaling our method for larger open source models is an area for future work.
+Method GQA Subset
+Answer Acc Stu/Tea Agree.
+Qwen Coder 2.5-0.5B (w/Aug) 43.2 70.5
+Qwen Coder 2.5-1.5B(w/ Aug) 42.5 65.8
+Qwen Coder 2.5-3B(w/ Aug) 42.4 67.6
+Table 10: When using $0.1\%$ of the training dataset, our method is more effective on smaller models,
+Appendix B Program Analysis
 Program Errors
-In Figure 3, we show the frequency of different error types for each method on the GQA dataset.
+In Figure 4, we show the frequency of different error types for Qwen-Coder-0.5B under few-shot model, non-augmented (joint) model and augmented (joint) model as well as the teacher model on the GQA dataset.
 Note that incorrect programs can fall into multiple categories but in this classification, each incorrect program was counted only once.
-There are very few execution errors and most errors come from either contradicting the question or not answering it.
-Many of the Llama generated programs returned yes/no even if the question asked for a different string.
-Augmentation largely reduces such errors but also generates new errors.
+Other than the few-shot model, there are no execution errors.
+The main sources of error are not answering question and API violation.
 Figure 4: The frequency of errors across the different categories for GQA program evaluation. Augmentation reduces the number of ‘Does Not Answer Question Mistakes.’
 Qualitative Analysis
-In Figure 4, we show 9 generated programs from 3 questions in the GQA dataset and 3 models: student model trained without augmentation, with augmentation and the auto-context teacher.
-Generated programs for existence questions about a single object and a single attribute like the one in the first column are almost always correct, even for the worse performing model, Llama-3.2-1B.
+In Figure 5, we show 9 generated programs from 3 questions in the GQA dataset and 3 models: CodeT5 student model trained without augmentation, CodeT5 with augmentation and the auto-context teacher.
+Generated programs for existence questions about a single object and a single attribute like the one in the first column are almost always correct, even for few-shot methods.
 In the second column, we see an example where the model trained without augmentation leaves out details mentioned in the question but the augmented model generates the correct program.
@@
 Simple comparison questions (left hand side) are almost always correct while questions with negations are almost always incorrect across the different methods.
-Appendix B Training and Model Details
+Appendix C Training and Model Details
 We used the following models for executing programs:
@@
 Program generation settings for GPT can be found in Table 11.
-Template-based and direct training hyper-parameters can be found in Table 11.
-For CodeT5, the most important hyper-parameters were the learning rate and LoRA dropout rate.
+Training hyper-parameters can be found in Table 12.
+For distilled models, the most important hyper-parameters were the learning rate and LoRA dropout rate.
 Training stopped when the training loss stopped decreasing.
-All experiments were run on a single 40gb A40 or 40gb A100. Time measurements were measured on an A40.
+All experiments were run on a single 40gb A40 or 40gb A100. Time measurements were taken on an A40.
 Setting Value
@@
 Max Output Tokens 256
-Table 11: Training and evaluation settings for CodeT5
-Appendix C Changes to ViperGPT API
+Table 12: Training and evaluation settings for student models. We use the same learning rate for all models.
+Appendix D Changes to ViperGPT API
 The following are major modifications made to the ViperGPT API Surís et al. (2023).
@@
 Table 13: Change in GQA test-dev accuracy using original ViperGPT API when not using a VLM when the program fails
-Appendix D Prompt
+Appendix E Prompt
 Instructions
@@
 - The last line of each program should be answer =
-Appendix E Variable Renamer
+Appendix F Variable Renamer
 class VariableRenamer(ast.NodeTransformer):
@@
 Given a generated program we call the function ‘replace_variables’ which uses and abstract-syntax tree to rename variables both in and outside different types of loops.
-Appendix F Variable Replacement
+Appendix G Variable Replacement
 Algorithm 1 Argument Replacement
@@
 Category Name Category Examples
-Color red, blue, green, yellow, purple, black, white, orange, pink, brown, gray, indigo, cyan magenta, tan, silver
-Activities running, walking, snowboarding, flying, splashing, tossing, riding, standing, hugging hanging, breaking, pulling, decorating, facing, preparing pouring, pointing, laughing
-Relation picking up, in front of behind, above, below next to, near, far away close , following, on top, beside, walking on, attached, left, right, diagonal
+Color
+red, blue, green, yellow,
+purple, black, white,
+orange, pink, brown,
+gray, indigo, cyan
+magenta, tan, silver
+Activities
+running, walking, snowboarding,
+flying, splashing, tossing,
+riding, standing, hugging
+hanging, breaking, pulling,
+decorating, facing, preparing
+pouring, pointing, laughing
+Relation
+picking up, in front of
+behind, above, below
+next to, near, far away
+close , following, on top,
+beside, walking on, attached,
+left, right, diagonal